# Research Challenges and Progress in the End-to-End V2X Cooperative Autonomous Driving Competition

### v1

@@ 6 Future Directions
The challenge results and observed limitations across both tracks highlight several key directions for advancing the field of cooperative autonomous driving under V2X settings. To bridge the gap between benchmark success and real-world deployment, future research should address the following critical aspects:
@@ 6.1 Realistic V2X Communication Modeling
Most current solutions assume ideal or simplified communication channels, with constant message delivery and no packet loss. However, real-world V2X systems often operate under non-deterministic network conditions, including variable latency, intermittent connectivity, and data dropouts due to interference or congestion. Future benchmarks and algorithms should incorporate communication-aware learning by simulating:
  • Packet loss models based on empirical wireless studies,
  • Delay-aware fusion mechanisms, where agents reason with stale or missing messages,
  • Redundancy-aware protocols, that prioritize critical information under constraints.
This would enable the design of robust agents that adapt their behavior not only based on perceptual uncertainty, but also on the reliability of the communication channel.
@@ 6.2 Bandwidth-Adaptive and Task-Aware Fusion
While sparse fusion strategies showed promise in this challenge, a crucial open question is how to dynamically adjust fusion strategies based on both bandwidth availability and task requirements. Future systems may benefit from:
  • Information-theoretic fusion policies, which select features with maximal utility for the downstream task,
  • Hierarchical encoding schemes, allowing agents to transmit coarse-to-fine updates depending on the link condition,
  • Task-specific prioritization, where planning-critical cues (e.g., dynamic agents near intersections) are communicated more aggressively than static context.
@@ 6.3 Generalization Across Heterogeneous Agents and Scenarios
Real-world autonomous systems will inevitably involve heterogeneous participants, including vehicles and infrastructure units with varying sensor types, fields of view, and computing capabilities. Robust fusion and planning under such heterogeneous conditions remain largely unsolved. Future research should explore:
  • Calibration-agnostic fusion frameworks, resilient to partial or inaccurate sensor alignment,
  • Meta-learning or domain adaptation techniques, to generalize across sensor configurations, cities, and deployment environments,
  • Scalable fusion topologies, that support dynamic participation (e.g., vehicles entering/leaving the scene).
Addressing this challenge will significantly improve the deployability of cooperative driving systems across different geographies and manufacturers.
@@ 6.4 Interpretability, Safety, and Standardization
For cooperative systems to be adopted in safety-critical applications such as autonomous driving, interpretability and verifiability become essential. Future work should focus on:
  • Transparent fusion architectures, that expose which agents and observations contributed to decisions,
  • Uncertainty quantification in cooperative predictions and plans,
  • Conformance to communication and safety standards, such as SAE J2735 or ETSI ITS-G5.
Moreover, establishing open benchmarks and evaluation protocols for interpretability and fault tolerance will further accelerate research translation into practice.
6.5 Community Building and Ecosystem Development
The field of V2X cooperative driving remains fragmented across perception, networking, and control communities. To build a coherent and impactful research direction, we encourage:
  • Continued development of open-source toolkits, such as UniV2X, that support full-stack experimentation,
  • Expansion of datasets to include adverse conditions (e.g., night, rain, sensor failure),
  • Organization of long-term multi-institutional benchmarks, fostering reproducibility, collaboration, and community-wide progress.
  • Increased support for V2X-specific challenges and competitions, as most existing benchmarks (see Table 1) remain focused on single-agent autonomy.
Through sustained infrastructure and shared challenge platforms, we can drive the maturation of V2X research from academic prototypes to robust, real-world systems.
@@ 7 Conclusion
This paper presented a comprehensive overview of the End-to-End Autonomous Driving through V2X Cooperation Challenge, held as part of the MEIS Workshop @ CVPR 2025. The challenge was designed to advance the state of cooperative autonomous driving by evaluating perception and planning systems under realistic multi-agent and communication-constrained conditions. It comprised two tracks—cooperative temporal perception and end-to-end planning—built on the open-source UniV2X framework and the V2X-Seq-SPD dataset.
Through participation from over 30 teams worldwide, the challenge revealed both significant progress and critical bottlenecks in V2X-enabled driving systems. Top solutions leveraged sparse, query-based fusion, modular architectures, and temporal reasoning to achieve strong results in both perception and planning. At the same time, key research challenges were identified in areas such as communication-aware fusion, robust planning under partial observability, and heterogeneous agent generalization.
The insights from this challenge underscore the importance of designing V2X systems that are not only accurate and efficient, but also adaptive, interpretable, and deployable under real-world constraints. By promoting open benchmarks, reproducible baselines, and community collaboration, this initiative aims to bridge the gap between academic research and practical deployment of cooperative autonomous driving technologies.
Future editions of the challenge will expand in scope and complexity, incorporating richer sensor setups, more realistic communication models, and diversified driving scenarios. We invite the broader research community to join in this effort to build safe, scalable, and intelligent multi-agent driving systems for the cities of tomorrow.
@@ Acknowledgement
The authors would like to express their sincere gratitude to all participating teams of the “End-to-End Autonomous Driving through V2X Cooperation” Challenge for their valuable contributions. In particular, we acknowledge the outstanding efforts of the team led by Ziyi Song and Dr. Sheng Zhou from Tsinghua University, as well as the team led by Dr. Ehsan Javanmardi from the University of Tokyo. We also gratefully acknowledge the sponsorship and support provided by the Multimedia Laboratory at the University of Hong Kong and Shanghai Songying Technology Co., Ltd. Furthermore, we sincerely acknowledge the support from the Wuxi Research Institute of Applied Technologies at Tsinghua University under Grant No. 20242001120.

### v2

@@ 6 Future Directions
The challenge results and observed limitations across both tracks reveal a clear gap between current benchmark performance and the requirements for real-world deployment of cooperative autonomous driving under V2X settings. To close this gap, future research should address the following interconnected directions, progressing from foundational communication challenges to system-level adoption.
@@ 6.1 Realistic V2X Communication Modeling
Most current solutions assume ideal or simplified communication channels with perfect message delivery [77, 83]. In practice, V2X networks suffer from variable latency, intermittent connectivity, and packet drops caused by interference or congestion [19, 59]. Future benchmarks and algorithms should embed communication-aware learning by:
  • Simulating packet loss models grounded in empirical wireless measurements [32],
  • Incorporating delay-aware fusion mechanisms that reason with stale or missing data [1, 76],
  • Designing redundancy-aware protocols or modules to prioritize safety-critical and planning-oriented information under bandwidth constraints [39].
These modules would enable robust agents that adapt to both perceptual uncertainty and communication reliability.
@@ 6.2 Bandwidth-Adaptive and Task-Aware Fusion
Fusion strategies must adapt to both bandwidth availability and task requirements to ensure reliable performance at scale [85]. However, most existing approaches remain static and fail to capture the dynamic trade-offs between efficiency and accuracy. Beyond sparse fusion, future feature fusion modules may benefit from:
  • Information-theoretic feature selection to maximize task utility per transmitted bit [52, 14],
  • Hierarchical encoding schemes for coarse-to-fine updates based on link conditions [46, 50],
  • Task-driven prioritization, sending safety-critical and planning-critical cues (e.g., nearby dynamic agents) more aggressively than static context.
@@ 6.3 Generalization Across Heterogeneous Agents and Scenarios
Real-world deployments will involve heterogeneous vehicles and infrastructure with varying sensors, fields of view, and computational capabilities [89, 65, 70]. Robust fusion and planning under such diversity can be supported by:
  • Calibration-agnostic fusion mechanisms or frameworks tolerant to partial or inaccurate alignment [27],
  • Meta-learning or domain adaptation methods to generalize across sensor setups, cities, and conditions [45, 43],
  • Scalable fusion topologies that support dynamic participation as agents enter or leave the scene [62].
Addressing these challenges will significantly improve the deployability of cooperative driving systems across different geographies and manufacturers.
6.4 Air-Ground Collaboration
The growth of the low-altitude economy [67] will introduce drones as additional sensing and communication agents for autonomous driving [30, 64, 35, 26]. By acting as “free viewpoints,” drones can enrich situational awareness in dense traffic and occluded intersections. At the same time, their deployment raises issues of sensor vibration, communication overhead, and limited endurance. Air-ground cooperative systems can be advanced by:
  • Addressing sensor vibration artifacts in drone-mounted perception, which can degrade image quality and downstream detection and planning performance [4],
  • Developing bandwidth-aware fusion algorithms for transmitting aerial data efficiently,
  • Coordinating multiple drones under battery and flight-time constraints for persistent coverage [13, 17].
Leveraging aerial viewpoints in coordination with ground vehicles could unlock richer situational awareness and enable safer, large-scale deployment of cooperative driving systems.
@@ 6.5 Interpretability, Safety, and Standardization
For cooperative systems to be adopted in safety-critical applications such as autonomous driving, interpretability, verifiability and standardization become essential [44, 87, 2]. Progress can be made through:
  • Transparent fusion architectures that expose the contribution of each agent and observation [82, 44],
  • Uncertainty quantification in cooperative perception and planning outputs [47, 7],
  • Conformance to communication and safety standards such as SAE J2735 [60] or ETSI ITS-G5 [63].
These directions are critical to ensuring that cooperative driving systems are not only performant but also trustworthy and certifiable for real-world use.
6.6 Language as Communication Medium
Recent advances in Large Vision-Language Models [42, 75, 68] have opened the possibility of using natural language as a medium for V2X communication [29, 28, 21, 51, 71, 16, 31, 72]. Language-based communication promises transparency, efficiency, and interoperability, but its deployment in safety-critical driving scenarios remains largely unexplored. Future research should address:
  • Developing structured and unambiguous protocols for language-based V2X exchanges to avoid ambiguity [28, 71],
  • Combining natural language with traditional feature-level or state-level communication in hybrid pipelines,
  • Studying robustness to multilingual, noisy, or adversarial language inputs in cooperative driving,
  • Exploring decision-level negotiation mechanisms that go beyond perception sharing [22].
Addressing these challenges would transform natural language from a promising idea into a practical medium for cooperative autonomous driving.
6.7 Community and Ecosystem Development
Progress in V2X cooperative driving will be accelerated by cohesive community efforts and shared infrastructure. Key steps include:
  • Continuing development of open-source toolkits, such as UniV2X, for full-stack experimentation,
  • Expanding datasets to cover adverse conditions (e.g., night, rain, sensor failures),
  • Establishing long-term multi-institutional benchmarks to ensure reproducibility and collaboration,
  • Organizing V2X-specific competitions to stimulate innovation beyond single-agent autonomy.
  • Considering broader ethical and policy implications of multi-agent cooperation, including inter-manufacturer trust and data governance among automotive original equipment manufacturers (OEMs), as well as data security and regulatory compliance.
Building such community resources and ecosystems will help translate academic advances into robust, deployable cooperative driving systems worldwide.
@@ 7 Conclusion
This paper has presented a comprehensive overview of the End-to-End Autonomous Driving through V2X Cooperation Challenge, organized as part of the MEIS Workshop @ CVPR 2025. The challenge was designed to advance the state of cooperative autonomous driving by rigorously evaluating perception and planning systems under realistic multi-agent and communication-constrained conditions. It comprised two tracks: cooperative temporal perception and end-to-end planning, built upon the open-source UniV2X framework [87] and the V2X-Seq-SPD dataset [86].
Participation from numerous teams highlighted both substantial progress and persistent limitations in V2X-enabled driving systems. The strongest submissions adopted sparse, query-based fusion, modular architectures, and temporal reasoning, achieving competitive results in both perception and planning tasks. At the same time, we summarize the critical open challenges, including communication-aware fusion, robust planning under partial observability, and generalization across heterogeneous agents.
These findings and insights from related research, emphasize that the development of cooperative driving systems must go beyond accuracy and efficiency, extending to adaptability, interpretability, and robustness in real-world deployments. The challenge has further underscored the importance of open benchmarks, reproducible baselines, and cross-community collaboration as essential drivers for translating academic innovation into deployable systems.
Looking forward, future editions of the challenge will broaden their scope to incorporate richer sensor modalities, more realistic and dynamic communication models, and increasingly diverse driving environments. By continuing to integrate technical advances with ethical, regulatory, and societal considerations, this initiative seeks to foster the design of safe, scalable, and intelligent multi-agent driving systems for the urban mobility ecosystems of tomorrow.
@@ Acknowledgement
The authors would like to express their sincere gratitude to all participating teams of our challenge for their valuable contributions. In particular, we acknowledge the outstanding efforts of the team led by Ziyi Song and Dr. Sheng Zhou from Tsinghua University, as well as the team led by Dr. Ehsan Javanmardi from the University of Tokyo. We would also like to thank Xiangbo Gao and Dr. Zhengzhong Tu from Texas A&M University for their valuable suggestions on future work. We also gratefully acknowledge the sponsorship provided by the Multimedia Laboratory at the University of Hong Kong and Shanghai Songying Technology Co., Ltd. Furthermore, we sincerely acknowledge the support from the Wuxi Research Institute of Applied Technologies at Tsinghua University under Grant No. 20242001120.